8) Natal Corp. has issued some $100 par preferred stock with a dividend of 10% of par value. The required rate of return on the preferred stock is 9%.

a) Compute the current value of the preferred stock.

b) If the preferred stock sells in the market at $115 per share, what is the expected rate of return from the preferred stock.

c) If Natal announces that it will redeem the preferred stock in 4 years at a redemption price of $140 per share, compute the rate of return from the preferred stock.
Answer #1

a.

Annual Dividend = 0.10(100) = $10

Current Value of Preferred Stock = 10/0.09 = $111.11

b.

Expected Return = 10/115 = 8.70%

c.

Rate of Return = 10/140 = 7.14%